Determine the partial pressures of ideal gases A, B, and C in a container with a total pressure of 3.5 atm. The number of moles for Gas A is twice that of Gas B and four times that of Gas C.


1
Expert's answer
2021-01-22T06:23:02-0500

A = 2x , B = x , C = 0.5x

A+B+C = 3.5atm

2x+x+0.5x = 3.5atm

3.5x = 3.5atm

x = 1 atm

A = 2x = 2×1 = 2atm

B = x = 1 atm

C = 0.5x = 0.5×1 = 0.5atm